# Scalable piecewise smoothing with BART

### v3

@@ 1.1 Motivation: Smooth NBA shot charts
Figures 1(a) and 1(b) show the locations of made and missed shots for two players, Stephen Curry and Ben Simmons, from the 2023-24 National Basketball Association season. One way to estimate shot charts is to divide the court into several small regions and to compute the proportion of shots each player makes in each region. Unfortunately, such an approach can produce sharp discontinuities. Consider region A in Figure 1(a): Curry made three of his six attempts in this region but made all five of his attempts in the region immediately to the right of A. The resulting sharp transition from 50% to 100% is highly undesirable as player’s shooting abilities are generally believed to vary in spatially smooth fashion (Yin et al.,, 2022, 2023). A further difficulty lies in predicting FGat locations where a player has not attempted a shot. As an extreme example, Ben Simmons did not attempt a single shot within or near the region B in Figure 1(b). The naïve binning-and-average procedure provides little principled guidance for predicting Simmons’ FGat these locations.
A Bayesian hierarchical model is a conceptually simple way to overcome these difficulties. For example, one can express each player’s probit-transformed FGas a linear combination of fixed functions of locations (e.g., a tensor product of splines) and specify an exchangeable prior for the player-specific coefficients. Such a model “borrows strength” across players so that every player’s estimated FGat a particular location is informed not only by their own data but also by the estimated FG%’s of all other players at that location. So, a simple hierarchical model would shrink Simmons’ FGin region B towards a league-wide average.
@@
Ultimately, FGdepends on complex interactions between location, player, and player characteristics like position, height, and weight. The cases of Curry and Simmons underscore the difficulties faced in trying to specify these interactions correctly in a parametric model. On this view, Bayesian Additive Regression Trees (BART; Chipman et al.,, 2010), which approximates regression functions using an ensemble of binary regression trees, offers a compelling alternative. With BART, users often obtain excellent predictive results and reasonably calibrated uncertainties “off-the-shelf” without needing to pre-specify the functional form of the regression function and with very little hyperparameter tuning.
Figure 1: Locations of made and missed shots for Stephen Curry (a) and Ben Simmons (b). BART-estimated FGas a function for location for Curry (c) and Simmons (d).
Because BART is based on axis-aligned regression trees, the BART-based shot charts in Figures 1(c) and 1(d) display characteristic horizontal and vertical striations. Since these trees are nothing more than piecewise constant step functions, BART-based shot charts will generally not be smooth and will instead contain sharp discontinuities. For instance, Curry’s and Simmons’ estimated FGjump from below 50% (blue shading) to above 50% (orange/red shading) and then back to below 50% within regions C, D, and E in the figures.
Several authors have proposed BART extensions based on regression trees that output smooth functions including linear models (Prado et al.,, 2021), splines (Low-Kam et al.,, 2015; Cao et al.,, 2025), and Gaussian Processes (GPs; Starling et al.,, 2020; Maia et al., 2024a, ). Given the ubiquity of splines and GPs in spatial statistics, the latter two types of extensions are attractive options for estimating shot charts. Unfortunately, existing implementations either (i) only permit smoothing in one dimension (Starling et al., (2020)’s tsBART and the spline models of Low-Kam et al., (2015) and Cao et al., (2025)) or (ii) scale poorly and cannot be fit to a single’s season of data consisting of about 190,000 shots (tsBART and Maia et al., 2024a ’s GP-BART).
@@ 1.2 Our Contribution
We develop a novel extension of BART, which we call ridgeBART, that uses trees that output linear combinations of ridge functions, which compose affine transformations of the inputs with a non-linearity. Whereas the original BART model approximates functions with piecewise constant step functions, ridgeBART approximates functions with piecewise continuous functions. We introduce a new Gibbs sampler whose complexity scales linearly with the size of the data, offering substantial runtime improvements over GP-based BART extensions. We further derive posterior contraction rates for the ridgeBART posterior in the setting where the true regression function lies in a piecewise anisotropic Hölder class. Briefly, this class contains functions that display different degrees of smoothness with respect to each covariate in different rectangular sub-regions of $[0,1]^{p}.$ As a special case of our main result, we show that the ridgeBART posterior contracts at almost the minimax optimal rate when the true regression function belongs to the isotropic Hölder class $\mathcal{C}^{\alpha}([0,1]^{p})$ for arbitrary $\alpha>0,$ losing only a $\log{(n)}$-factor. An R implementation of ridgeBART is available at https://github.com/ryanyee3/ridgeBART.
The rest of the paper is structured as follows: in Section 2, we review the basics of the original BART model and its GP-based extensions. Then, we formally introduce ridgeBART and derive our Gibbs sampling algorithm in Section 3 before studying its theoretical properties in Section 4. In Section 5, we assess ridgeBART’s empirical performance using several synthetic datasets and use it to produce NBA shot charts. We conclude in Section 6 by discussing potential methodological and theoretical extensions.
@@ 2 A review of BART and its GP-based extensions
Chipman et al., (2010) introduced BART in the context of the standard nonparametric regression problem: given $n$ observations from the model $y\sim\mathcal{N}\left(f(\bm{x}),\sigma^{2}\right)$ with $\bm{x}\in[0,1]^{p},$ BART expresses $f$ as a sum of a large number of binary regression trees.
A regression tree is a pair $(\mathcal{T},\mathcal{M})$ of (i) a finite, rooted binary decision tree $\mathcal{T}$ consisting of $\mathcal{L}(\mathcal{T})$ terminal or leaf nodes and several non-terminal or decision nodes; and (ii) a collection of jumps =$\mathcal{M}=\{\mu_{1},\ldots,\mu_{\mathcal{L}(\mathcal{T})}\}$, one for each leaf node in $\mathcal{T}.$ The decision tree partitions $[0,1]^{p}$ into several axis-parallel boxes, one for each leaf, and we can associate each $\bm{x}\in[0,1]^{p}$ with a single leaf denoted $\ell(\bm{x};\mathcal{T}).$ By further associating each leaf $\ell$ with a scalar jump $\mu_{\ell},$ the regression tree $(\mathcal{T},\mathcal{M})$ represents a piecewise constant step function. We let $g(\bm{x};\mathcal{T},\mathcal{M})=\mu_{\ell(\bm{x};\mathcal{T})}$ denote the evaluation function, which returns the jump associated to $\bm{x}$’s leaf.
@@ 2.1 The original BART model
BART introduces an ensemble =$\mathcal{E}=\{(\mathcal{T}_{1},\mathcal{M}_{1}),\ldots,(\mathcal{T}_{M},\mathcal{M}_{M})\}$ of $M$ regression trees so that $f(\bm{x})=\sum_{m=1}^{M}{g(\bm{x};\mathcal{T}_{m},\mathcal{M}_{m})}.$ Fitting a BART model involves specifying a prior over $(\mathcal{E},\sigma)$ and computing summaries of the posterior distribution of $(\mathcal{E},\sigma)|\bm{y}$ using Markov chain Monte Carlo (MCMC).
The BART prior. Chipman et al., (2010) specified a conjugate inverse gamma prior for $\sigma^{2}$ and independent and identical priors on the regression trees. We describe that prior implicitly by explaining how to simulate a random prior sample. First, we use a branching process to generate the graphical structure of $\mathcal{T}$ (i.e., the arrangement of decision and leaf nodes). In the branching process, the probability that a node at depth $d$ is non-terminal is $0.95(1+d)^{-2},$ which ensures that $\mathcal{T}$ is almost surely finite. Then, we draw decision rules for each non-terminal node conditionally given the rules at all of its ancestors in the tree, which ensures each leaf of $\mathcal{T}$ corresponds to a non-empty subset of $[0,1]^{p}$ with prior probability one. Finally, conditionally on $\mathcal{T},$ we independently draw jumps $\mu_{\ell}\sim\mathcal{N}\left(0,\tau^{2}\right).$ Chipman et al., (2010) suggested default values for $\tau$ and an inverse Gamma prior for $\sigma^{2}$ that have proven highly effective across many applications.
Sampling from the BART posterior. Chipman et al., (2010) deployed a Gibbs sampler that iterated between sampling each of $\mathcal{E}$ and $\sigma^{2}$ conditionally given the other. Sampling $\sigma^{2}|\bm{y},\mathcal{E}$ involves a straightforward normal-inverse Gamma conjugate update. For the more involved conditional update of $\mathcal{E},$ Chipman et al., (2010) used a Metropolis-within-Gibbs strategy that sequentially samples one regression tree at a time, keeping all others fixed. Rather than updating the decision tree and jumps simultaneously, they first drew the decision tree marginally and then conditionally sampled the jumps.
@@
The marginal decision tree density and the conditional jump distributions are available in closed-form and can be computed in $O(n)$ time. To describe these calculations, suppose we are updating the $m$-th regression tree $(\mathcal{T}_{m},\mathcal{M}_{m})$ while leaving the other $M-1$ regression trees, which we denote by $\mathcal{E}^{(-m)},$ fixed. Let $r_{i}$ be the $i$-th partial residual based on the fits of the other trees in $\mathcal{E}^{(-m)}$ with $r_{i}=y_{i}-\sum_{m^{\prime}\neq m}{g(\bm{x}_{i};\mathcal{T}_{m^{\prime}},\mathcal{M}_{m^{\prime}})}.$ Additionally, for each leaf $\ell$ of an arbitrary decision tree $\mathcal{T},$ let $\mathcal{I}(\ell;\mathcal{T})=\{i:\ell(\bm{x}_{i};T)=\ell\}$ be the set of indices for those observations associated with leaf $\ell$ and let $n_{\ell}=\lvert\mathcal{I}(\ell;\mathcal{T})\rvert$ count the number of observations associated with leaf $\ell.$ Finally, let $\bm{r}_{\ell}=(r_{i}:i\in\mathcal{I}(\ell;\mathcal{T}))$ be the vector of partial residuals corresponding to observations in leaf $\ell.$ The conditional posterior distribution of $(\mathcal{T}_{m},\mathcal{M}_{m})$ factorizes over the leafs of the tree:
@@
where $P_{\ell}=\sigma^{-2}n_{\ell}+\tau^{-2}$ and $\Theta_{\ell}=\sigma^{-2}\bm{r}^{\top}_{\ell}\bm{1}_{n_{\ell}}.$ Equation 1 further reveals that the jumps $\mu_{\ell}\in\mathcal{M}_{m}$ are conditionally independent with $\mu_{\ell}|\mathcal{T}_{m},\bm{y},\mathcal{E}^{(-m)},\sigma^{2}\sim\mathcal{N}\left(P_{\ell}^{-1}\Theta_{\ell},P_{\ell}^{-1}\right).$ Thus, the main computation involved in updating each regression tree is summing the partial residuals in each leaf $\ell,$ which takes $O(n_{\ell})$ time. Summing over each leaf, we conclude that each individual regression tree updates requires $O(n)$ time.
@@ 2.2 Updating smoother regression trees
Unlike the original BART model, each regression tree update in Starling et al., (2020)’s and Maia et al., 2024a ’s extensions cannot be done in linear time. Indeed, each update involves computing the Cholesky factorizations of $P_{\ell}$ and $K_{\ell},$ from which both the determinant and inverse can be computed. Consequently, each regression tree update requires $O(\sum_{\ell}{n_{\ell}^{3}})$ time, which can be computationally prohibitive when $n$ is large.
@@ 3 Introducing ridgeBART
By utilizing scalar jumps, the original BART model admits very fast regression tree updates but is limited to representing discontinuous functions. GP-based extensions of BART, on the other hand, represent a much richer set of functions at the expense of much slower computation. To achieve a middle ground maintaining the representational flexibility of treed-GP ensembles while admitting linear-time tree updates, we express the functional jump $\mu_{\ell}(\bm{z})$ as a linear combination of $D$ ridge functions
@@
Motivation for Equation 4. Before proceeding, we pause to motivate the form of our tree outputs. Suppose that $k:[0,1]^{q}\times[0,1]^{q}\rightarrow\mathbb{R}$ is a real-valued stationary kernel and that $h\sim\textrm{GP}(0,k).$ Given observations $h(\bm{z}_{1}),\ldots,h(\bm{z}_{n}),$ sampling from the posterior distribution of $h(\bm{z})$ involves computing the Cholesky factorization of the $n\times n$ kernel matrix with entries $k(\bm{z}_{i},\bm{z}_{i^{\prime}})$ (see, e.g., Rasmussen and Williams,, 2006, Equation (2.19) and Appendix A2). Exactly computing this factorization requires $O(n^{3})$ time, which becomes computationally prohibitive when $n$ is large. Rahimi and Recht, (2007) proposed the approximation $k(\bm{z},\bm{z}^{\prime})=2D^{-1}\sum_{d=1}^{D}{\cos\left(\omega_{d}^{\top}\bm{z}+b_{d}\right)\cos\left(\omega_{d}^{\top}\bm{z}^{\prime}+b_{d}\right)},$ where $b_{d}\sim\textrm{Uniform}\left(0,2\pi\right)$ and the $\omega_{d}$’s are drawn i.i.d. from $k$’s spectral measure $\mathcal{P}_{\omega},$ which satisfies $k(\bm{z},\bm{z}^{\prime})=\mathbb{E}_{\omega\sim\mathcal{P}_{\omega}}[e^{i\omega^{\top}(\bm{z}-\bm{z}^{\prime})}]$ [^2] The existence of $\mathcal{P}_{\omega}$ is guaranteed by Bochner’s theorem.. Essentially, their result allows us to approximate the function $h$ as a linear combination of $D$ randomly constructed cosine features. Computing the posterior of the weights, which induces an approximate posterior over $h(\bm{z}),$ requires only $O(nD^{2})$ time, a considerable speedup over the $O(n^{3})$ time needed for exact GP computation.
Li et al., (2023) built regression trees ensembles with random Fourier feature expansions in each leaf. While similar to that work, our proposed model, theoretical results, and software implementation are considerably more general, allowing for a wide range of popular activation functions $\varphi(\cdot)$ like the hyperbolic tangent $\tanh(\cdot),$ the sigmoid $[1+e^{-\cdot}]^{-1},$ and the rectified linear unit (ReLU) $\max\{0,\cdot\}.$ By associating each tree leaf with an expansion like Equation 4, we are effectively building an ensemble of treed-neural networks.

### v4

@@ 1.1 Motivation: Smooth NBA shot charts
Figures 1(a) and 1(b) show the locations of made and missed shots for two players, Stephen Curry and Ben Simmons, from the 2023-24 National Basketball Association season. One way to estimate shot charts is to divide the court into several small regions and to compute the proportion of shots each player makes in each region. Unfortunately, such an approach can produce sharp discontinuities. Consider region A in Figure 1(a): Curry made three of his six attempts in this region but made all five of his attempts in the areas immediately to the right of region A. The resulting sharp transition from 50% to 100% is highly undesirable as players’ shooting abilities vary in a spatially smooth fashion (Yin et al.,, 2022, 2023). A further difficulty lies in predicting FG% at locations where a player has not attempted a shot. As an extreme example, Ben Simmons did not attempt a single shot within or near the region B in Figure 1(b). Simple binning and averaging precludes principled prediction of Simmons’ FG% at these locations.
Bayesian hierarchical models are a conceptually simple way to overcome these difficulties. For example, one can express each player’s probit-transformed FG% as a linear combination of fixed functions of locations (e.g., a tensor product of splines) and specify an exchangeable prior for the player-specific coefficients. Such a model “borrows strength” across players so that every player’s estimated FG% at a particular location is informed not only by their own data but also by the estimated FG%’s of all other players at that location.
@@
Ultimately, FG% depends on complex interactions between location, player, and player characteristics like position, height, and weight. The cases of Curry and Simmons underscore the difficulties faced in trying to specify these interactions correctly in a parametric model. On this view, Bayesian Additive Regression Trees (BART; Chipman et al.,, 2010), which approximates regression functions using an ensemble of binary regression trees, offers a compelling alternative. With BART, users often obtain excellent predictive results and reasonably calibrated uncertainties “off-the-shelf” without needing to pre-specify the functional form of the regression function and with very little hyperparameter tuning.
Figure 1: Locations of made and missed shots for Stephen Curry (a) and Ben Simmons (b). BART-estimated FG% as a function for location for Curry (c) and Simmons (d).
Figures 1(c) and 1(d) show the estimated shot charts for Curry and Simmons obtained from a BART model whose trees split on continuous features like shot location, player height, and player weight and categorical features like player name and player position. Since these trees are nothing more than piecewise constant step functions, these shot display contain sharp discontinuities in location. For instance, Curry’s and Simmons’ estimated FG% jump from below 50% (blue shading) to above 50% (orange/red shading) and then back to below 50% within regions C, D, and E in the figures.
Several authors have proposed BART extensions based on regression trees that output smooth functions including linear models (Prado et al.,, 2021), splines (Low-Kam et al.,, 2015; Cao et al.,, 2025), and Gaussian Processes (GPs; Starling et al.,, 2020; Maia et al., 2024a, ). Given the ubiquity of splines and GPs in spatial statistics, the latter two types of extensions are attractive options for estimating shot charts. Unfortunately, existing implementations either (i) only smooth in one dimension (Starling et al., (2020)’s tsBART and Low-Kam et al., (2015)’s and Cao et al., (2025)’s treed-spline models) or (ii) scale poorly and cannot be fit to a single’s season of data consisting around 190,000 shots (tsBART and Maia et al., 2024a ’s GP-BART).
@@ 1.2 Our Contribution
We develop a novel extension of BART, which we call ridgeBART, that uses trees that output linear combinations of ridge functions, which compose affine transformations of the inputs with a non-linearity. Whereas the original BART model approximates functions with piecewise constant step functions, ridgeBART approximates functions with piecewise continuous functions. We introduce a new Gibbs sampler whose complexity scales linearly with the size of the data, offering substantial runtime improvements over GP-based BART extensions. We further derive posterior contraction rates for the ridgeBART posterior in the setting where the true regression function lies in a piecewise anisotropic Hölder class. Briefly, this class contains functions that display different degrees of smoothness with respect to each covariate in different rectangular sub-regions of $0,1]^{p}.$ As a special case of our main result, we show that the ridgeBART posterior contracts at almost the minimax optimal rate when the true regression function belongs to the isotropic Hölder class $\mathcal{C}^{\alpha}([0,1]^{p})$ for arbitrary $\alpha>0,$ losing only a $\log{(n)}$-factor. An R implementation of ridgeBART is available at https://github.com/ryanyee3/ridgeBART.
In Section 2, we review the basics of the original BART model and its GP-based extensions. Then, we formally introduce ridgeBART and derive our Gibbs sampling algorithm in Section 3 before studying its theoretical properties in Section 4. In Section 5, we assess ridgeBART’s empirical performance using several synthetic datasets and use it to produce NBA shot charts. We conclude in Section 6 by discussing potential methodological and theoretical extensions.
@@ 2 A review of BART and its GP-based extensions
Given $n$ observations from the model $y\sim\mathcal{N}\left(f(\bm{x}),\sigma^{2}\right)$ with $\bm{x}\in[0,1]^{p},$ BART expresses $f$ as a sum of a large number of binary regression trees.
A regression tree is a pair $(\mathcal{T},\mathcal{M})$ of (i) a finite, rooted binary decision tree $\mathcal{T}$ consisting of $\mathcal{L}(\mathcal{T})$ terminal or leaf nodes and several non-terminal or decision nodes; and (ii) a collection of jumps $\mathcal{M}=\{\mu_{1},\ldots,\mu_{\mathcal{L}(\mathcal{T})}\}$, one for each leaf node in $\mathcal{T}.$ The decision tree partitions $0,1]^{p}$ into several axis-parallel boxes, one for each leaf, and we can associate each $\bm{x}\in[0,1]^{p}$ with a single leaf denoted $\ell(\bm{x};\mathcal{T}).$ By further associating each leaf $\ell$ with a scalar jump $\mu_{\ell},$ the regression tree $(\mathcal{T},\mathcal{M})$ represents a piecewise constant step function. We let $g(\bm{x};\mathcal{T},\mathcal{M})=\mu_{\ell(\bm{x};\mathcal{T})}$ denote the evaluation function, which returns the jump associated to $\bm{x}$’s leaf.
@@ 2.1 The original BART model
BART introduces an ensemble $\mathcal{E}=\{(\mathcal{T}_{1},\mathcal{M}_{1}),\ldots,(\mathcal{T}_{M},\mathcal{M}_{M})\}$ of $M$ regression trees so that $f(\bm{x})=\sum_{m=1}^{M}{g(\bm{x};\mathcal{T}_{m},\mathcal{M}_{m})}.$ Fitting a BART model involves specifying a prior over $(\mathcal{E},\sigma)$ and computing summaries of the posterior distribution of $(\mathcal{E},\sigma)|\bm{y}$ using Markov chain Monte Carlo (MCMC).
The BART prior. Chipman et al., (2010) specified a conjugate inverse gamma prior for $\sigma^{2}$ and independent and identical priors on the regression trees. We describe that prior implicitly by explaining how to simulate a random prior sample. First, we use a branching process to generate the graphical structure of $\mathcal{T}$ (i.e., the arrangement of decision and leaf nodes). In the branching process, the probability that a node at depth $d$ is non-terminal is $0.95(1+d)^{-2},$ which ensures that $\mathcal{T}$ is almost surely finite. Then, we draw decision rules for each non-terminal node conditionally given the rules at all of its ancestors in the tree, which ensures each leaf of $\mathcal{T}$ corresponds to a non-empty subset of $0,1]^{p}$ with prior probability one. Finally, conditionally on $\mathcal{T},$ we independently draw jumps $\mu_{\ell}\sim\mathcal{N}\left(0,\tau^{2}\right).$ Chipman et al., (2010) suggested default values for $\tau$ and an inverse Gamma prior for $\sigma^{2}$ that have proven highly effective across many applications.
Sampling from the BART posterior. Chipman et al., (2010) deployed a Gibbs sampler that iterated between conditionally sampling $\mathcal{E}$ and $\sigma^{2}.$ Sampling $\sigma^{2}|\bm{y},\mathcal{E}$ involves a normal-inverse Gamma conjugate update. For the more involved conditional update of $\mathcal{E},$ they used a Metropolis-within-Gibbs strategy that sequentially samples one regression tree at a time, keeping all others fixed, by first drawing a new decision tree marginally before conditionally sampling new jumps.
@@
The marginal decision tree density and the conditional jump distributions are available in closed-form and can be computed in $O(n)$ time. To describe these calculations, suppose we are updating the $m$-th regression tree $(\mathcal{T}_{m},\mathcal{M}_{m})$ while leaving the other M1M-1 regression trees, which we denote by $\mathcal{E}^{(-m)},$ fixed. Let $r_{i}$ be the $i$-th partial residual based on the fits of the other trees in $\mathcal{E}^{(-m)}$ with $r_{i}=y_{i}-\sum_{m^{\prime}\neq m}{g(\bm{x}_{i};\mathcal{T}_{m^{\prime}},\mathcal{M}_{m^{\prime}})}.$ Additionally, for each leaf $\ell$ of an arbitrary decision tree $\mathcal{T},$ let $\mathcal{I}(\ell;\mathcal{T})=\{i:\ell(\bm{x}_{i};T)=\ell\}$ be the set of indices for those observations associated with leaf $\ell$ and let $n_{\ell}=\lvert\mathcal{I}(\ell;\mathcal{T})\rvert$ count the number of observations associated with leaf $\ell.$ Finally, let $\bm{r}_{\ell}=(r_{i}:i\in\mathcal{I}(\ell;\mathcal{T}))$ be the vector of partial residuals corresponding to observations in leaf $\ell.$ The conditional posterior distribution of $(\mathcal{T}_{m},\mathcal{M}_{m})$ factorizes over the leafs of the tree:
@@
where $P_{\ell}=\sigma^{-2}n_{\ell}+\tau^{-2}$ and $\Theta_{\ell}=\sigma^{-2}\bm{r}^{\top}_{\ell}\bm{1}_{n_{\ell}}.$ Equation 1 further reveals that the jumps $\mu_{\ell}\in\mathcal{M}_{m}$ are conditionally independent with $\mu_{\ell}|\mathcal{T}_{m},\bm{y},\mathcal{E}^{(-m)},\sigma^{2}\sim\mathcal{N}\left(P_{\ell}^{-1}\Theta_{\ell},P_{\ell}^{-1}\right).$ Since summing the partial residuals in each leaf $\ell$ takes $O(n_{\ell})$ time, we conclude that each individual regression tree update requires $O(n)$ time.
@@ 2.2 Updating smoother regression trees
Unlike the original BART model, each regression tree update in Starling et al., (2020)’s and Maia et al., 2024a ’s extensions cannot be done in linear time. Indeed, each update involves computing the Cholesky factorizations of $P_{\ell}$ and $K_{\ell}.$ Consequently, each regression tree update requires $O(\sum_{\ell}{n_{\ell}^{3}})$ time, which can be computationally prohibitive when $n$ is large.
@@ 3 Introducing ridgeBART
The original BART model admits very fast regression tree updates but is limited to representing discontinuous functions. GP-based extensions of BART, on the other hand, represent a much richer set of functions at the expense of much slower computation. To achieve a middle ground maintaining the representational flexibility of treed-GP ensembles while admitting linear-time tree updates, we express the functional jump $\mu_{\ell}(\bm{z})$ as a linear combination of $D$ ridge functions
@@
Motivation for Equation 4. Before proceeding, we pause to motivate the form of our tree outputs. Suppose that $k:[0,1]^{q}\times[0,1]^{q}\rightarrow\mathbb{R}$ is a real-valued stationary kernel and that $h\sim\textrm{GP}(0,k).$ Given observations $h(\bm{z}_{1}),\ldots,h(\bm{z}_{n}),$ sampling from the posterior distribution of $h(\bm{z})$ involves computing the Cholesky factorization of the $n\times n$ kernel matrix with entries $k(\bm{z}_{i},\bm{z}_{i^{\prime}})$ (see, e.g., Rasmussen and Williams,, 2006, Equation (2.19) and Appendix A2). Exactly computing this factorization requires $O(n^{3})$ time, which becomes computationally prohibitive when $n$ is large. Rahimi and Recht, (2007) proposed the approximation $k(\bm{z},\bm{z}^{\prime})=2D^{-1}\sum_{d=1}^{D}{\cos\left(\omega_{d}^{\top}\bm{z}+b_{d}\right)\cos\left(\omega_{d}^{\top}\bm{z}^{\prime}+b_{d}\right)},$ where $b_{d}\sim\textrm{Uniform}\left(0,2\pi\right)$ and the $\omega_{d}$’s are drawn i.i.d. from $k$’s spectral measure $\mathcal{P}_{\omega},$ which satisfies $k(\bm{z},\bm{z}^{\prime})=\mathbb{E}_{\omega\sim\mathcal{P}_{\omega}}[e^{i\omega^{\top}(\bm{z}-\bm{z}^{\prime})}]$ [^2] The existence of $\mathcal{P}_{\omega}$ is guaranteed by Bochner’s theorem.. Essentially, their result allows us to approximate $h$ with a linear combination of $D$ randomly constructed cosine features. Computing the posterior of the weights, which induces an approximate posterior over $h(\bm{z}),$ requires only $O(nD^{2})$ time, a considerable speedup over the $O(n^{3})$ time needed for exact GP computation.
Li et al., (2023) built regression trees ensembles with random Fourier feature expansions in each leaf. While similar to that work, our proposed model, theoretical results, and software implementation are considerably more general, allowing for a wide range of popular activation functions $\varphi(\cdot)$ like the hyperbolic tangent $\tanh(\cdot),$ the sigmoid $1+e^{-\cdot}]^{-1},$ and the rectified linear unit (ReLU) $\max\{0,\cdot\}.$ By associating each tree leaf with an expansion like Equation 4, we are effectively building an ensemble of treed-neural networks.